CG Beadix has daily average check collections of $101289, and it takes the firm 6 days before it can completely process those checks. An automated lockbox system that costs $303465 a year would reduce the processing time to 3 days. Beadix's opportunity cost of short-term funds is 12%. What is the net pretax benefit of the change? Round to the nearest dollar. Make sure to include a negative sign if it is negative.
Answer #1

AMount of funds released = Daily check amount*Days float reduced

= 101289*(6-3)  = 303867

Opportunity cost of funds (Benefit) = 12%*303867 = 36464

Net pretax benefit = Benefit from the system- cost

= 36464-303465

= - 267001